jhonyy9 (jhonyy9):

can we make this to result for every natural numbers ? 12345679*9 ------------ 111111111

jhonyy9 (jhonyy9):

this in case of result with 2 24691358*9 --------------- 222222222

jhonyy9 (jhonyy9):

so how will be for result of 3 ?

jhonyy9 (jhonyy9):

@agent0smith @welshfella @sweetburger @CandyCove

jhonyy9 (jhonyy9):

37*9 -------- 333 - just 3 terms of 3 can be resulted ? why ? can you explain it please ?

jhonyy9 (jhonyy9):

@ganeshie8 @TheSmartOne @mathmate @mathmath333 @mathmale

jhonyy9 (jhonyy9):

@Kainui

jhonyy9 (jhonyy9):

49382716*9 ------------- 444444444 1728395*9 -------------- 15555555 so not is right - or not possible ? 74*9 --------------- 666 in case of 6 these are all ? 86419753*9 -------------- 777777777 98765432*9 ---------------- 888888888

jhonyy9 (jhonyy9):

@pooja195 please your opinion ? ty

OpenStudy (kainui):

37037037 * 9 = 333333333 Not sure what you mean, that seems to work out ^

jhonyy9 (jhonyy9):

yes Kain - this is right for 3 - thank you can you make it for 6 too ?

jhonyy9 (jhonyy9):

and what or how will be for the case of 9 ?

OpenStudy (mathmate):

12345679*9=11111111 12345679*18=222222222 ... (12345679*9)*9=12345679*81=99999999 Actually they all come from 37. 37037037/3=12345679 so (37037037*9)*1=111111111 (37037037*9)*2=222222222 .... (37037037*9)*9=999999999 Now you can make it as long as you want: 37037037037037037/3=12345679012345679 (37037037037037037*9)=(12345679012345679*9)*1=111111111111111111 (37037037037037037*9)*2=(12345679012345679*9)*2=222222222222222222 ...

jhonyy9 (jhonyy9):

thank you @mathmate this is GREAT - you are the BEST !!!

jhonyy9 (jhonyy9):

@mathmath333 thank you to you too this website what is very interesting

OpenStudy (mathmath333):

np

OpenStudy (mathmate):

You're welcome! :)

jhonyy9 (jhonyy9):

@Directrix look please this too and wait your opinion about - ty